OpenStudy (anonymous):

How do you simplify this? (2x^2 * x)^3 I used * for multiplication

OpenStudy (johnweldon1993):

Basically do the FOIL method twice because you have (2x^2 * x)(2x^2 * x)(2x^2 * x) You just gotta multiply it out

OpenStudy (anonymous):

how do you multiply 2x^2 by x?

OpenStudy (zehanz):

\(2x^2 \cdot x=2x^3\). Now calculate \((2x^3)^3\) @johnweldon1993: I think you made a mistake about the FOIL method.

OpenStudy (johnweldon1993):

@ZeHanz thank you for pointing that out....FOIL wont work here sorry about that @stabar

OpenStudy (zehanz):

@stabar : remember: powers are just shorthand for repeated multiplication: \(a \cdot a=a^2\)

OpenStudy (zehanz):

So \(2x^2 \cdot x=2 \cdot x\cdot x\cdot x=2x^3\)

OpenStudy (anonymous):

right I get that...just not how I multipl that by x when i dont have a number for x

OpenStudy (zehanz):

You do not have to know how much x is. You cannot work out how much x*x is. You only can shorten it a little by writing it as x².

OpenStudy (anonymous):

ok so then 2x^2 * x would be 2x^3

OpenStudy (johnweldon1993):

Right...remember you can multiply any numbers together...remember that x....can also be written as 1x^1 And when multiplying eponents...you add them so yes 2x² * x WOULD be 2x^3

OpenStudy (anonymous):

so now all I have to do is work this out right?: (2x^3)(2x^3)(2x^3)

OpenStudy (zehanz):

Now the last step: \((2x^3)^3=2x^3 \cdot 2x^3 \cdot 2x^3=2 \cdot 2 \cdot 2 \cdot x^3 \cdot x^3 \cdot x^3=...\)

OpenStudy (zehanz):

You can take the 2's apart and also the factors x³.

OpenStudy (anonymous):

So I got 8x^9 is that right?

OpenStudy (zehanz):

Yes! You got it!

OpenStudy (anonymous):

Thank you

OpenStudy (zehanz):

YW!